PQRS is a rectangle in which PQ=9 cm and PS=6 cm.Tis a point on PQ such that PT=7 cm and RV is the perpendiculer from R to ST. Calculate ST and RV?
Answer by RAY100(1637)

need to draw sketch as defined above, rough is ok
to calc ST, we just have to recognize that triangle STP is a rt triangle, with PS=6, and PT=7 both given. using pythagorous st^2 = 6^2+7^2, and ST = 9.2195
to find RV, we need to recognize that PTS is similar to VSR ( note order of sides)
This follows from both have a rt angle(given), and angle PTS =angle TSR from AIA
sides are proportional in similar triangle, so being very careful to compare proportional sides
ST/RS=PS/RV
9.2/9=6/RV
(6*9)/(9.2)=RV
5.857=RV